OpenStudy (alprincenofl):

Use the Euclidean algorithm to find gcd (34, 21).

OpenStudy (alprincenofl):

Is this answer correct? We have 34 = 1•21+13, 21=1•13+8, 13=1•8+5, 8=1•5+3, 5=1•3+2, 3=1•2+1,2=2•1 Hence gcd(34,21) = 1.

OpenStudy (anonymous):

that looks good

OpenStudy (anonymous):

this one is annoying since you have to go all the way down

OpenStudy (alprincenofl):

Yes it is, that's why I asked to make sure :D thank you